OpenStudy (darkbeautystar):

Will Fan and Medal!! Given that squares X Y and Z are all similar, Square X has an area of 16. The reduced ratio between the area of square X and the area of Square Y is 1:3. The reduced ratio between the area of square Y and the area of square Z is 2:5. What is the reduced ratio between Squares X and Z?

OpenStudy (darkbeautystar):

@Luigi0210 -- Would you help me with this one question?

OpenStudy (darkbeautystar):

@mayankdevnani

OpenStudy (darkbeautystar):

@Awolflover1

OpenStudy (darkbeautystar):

BRB

OpenStudy (darkbeautystar):

Ok im back

OpenStudy (mayankdevnani):

where do you stuck ?

OpenStudy (darkbeautystar):

Well I am not exactly sure how to do this!

OpenStudy (mayankdevnani):

okay

OpenStudy (mayankdevnani):

\[\large \bf area~of~\boxed{x}=16sq.units\]

OpenStudy (mayankdevnani):

\[\large \bf \frac{area~\boxed{X}}{area~\boxed{Y}}=\frac{1}{3}=\frac{16}{area~\boxed{Y}}\]

OpenStudy (mayankdevnani):

find area of square Y?

OpenStudy (darkbeautystar):

Are you asking me if you have to do that?

OpenStudy (mayankdevnani):

find area of square Y

OpenStudy (mayankdevnani):

I can give you HINT only

OpenStudy (darkbeautystar):

Hangon

OpenStudy (mayankdevnani):

okay

OpenStudy (darkbeautystar):

I am not sure. :(

OpenStudy (darkbeautystar):

@mayankdevnani

OpenStudy (mayankdevnani):

tell me your answer

OpenStudy (darkbeautystar):

I didn't figure it out.... I want you to show me how to do it with out telling me the answer!

OpenStudy (darkbeautystar):

@mayankdevnani

OpenStudy (mayankdevnani):

cross multiply first

OpenStudy (darkbeautystar):

Cross multiply what?

OpenStudy (mayankdevnani):